A piece of metal has a mass of 200 g in air and 170 g when immersed in water. Calculate its volume and density. [Density of water is 1000 kgm-3. Take g=9.8 ms-2 ]

1
Expert's answer
2022-04-07T13:09:57-0400

The weight of a piece of metal in water

"W_{water}=W_{air}-g\rho_{water}V"

Hence, the volume of an object

"V=\frac{W_{air}-W_{water}}{g\rho_{water}}=\frac{0.20-0.17}{1000}=3*10^{-5}\:\rm m^3"

The density of object

"\rho=\frac{m}{V}=\frac{0.20\:\rm kg}{3*10^{-5}\:\rm m^3}=6700\:\rm kg/m^3"